# Observer Design for Networked Linear Systems with Fast and Slow Dynamics under Measurement Noise

### v3

@@ Abstract
thanks: This work was supported by the Australian Research Council under the Discovery Project DP200101303, the France Australia collaboration project IRP-ARS CNRS and the ANR COMMITS ANR-23-CE25-0005, and by ANID through the FONDECYT Initiation grant 11250715. E-mail corresponding author: weixuanw@student.unimelb.edu.au. A full version of this paper is available on https://arxiv.org/abs/2511.16469, see wang2025observer.
@@ 1 Introduction
This work provides two main contributions. First, it introduces a systematic observer design framework for SPNCS with linear plants via LMI. In particular, the proposed observer explicitly incorporates key network-induced effects, including sampling and scheduling protocols. By doing so, it enables the derivation of observer gains that make the observer robust against network-induced errors and thereby maximize the MATI estimates for both slow- and fast-varying signals. This contribution provides a unified and computationally tractable method for observer synthesis in multi-time-scale NCS, where network constraints play a critical role. Second, the paper establishes a robustness analysis framework for observer design in SPNCS with linear plants subject to measurement noise and plant input. In particular, we establish a DISS-like property for the estimation error, showing that the ultimate bound scales with the magnitudes of the measurement noises, as they are sampled and transmitted, and with the time derivative of the control inputs. These results not only quantify the sensitivity of the closed-loop system to different types of disturbances but also demonstrate how robustness relates to the system parameters, the protocol and MATI estimates.
@@ 3.1 Change of coordinates
Define the estimation errors on the slow and fast states as $\delta_{x}\coloneqq x_{o}-x_{p}$ and $\delta_{z}\coloneqq z_{o}-z_{p}$, respectively. Let $\hat{v}\coloneqq(\hat{v}_{1},\hat{v}_{2})$ and $v\coloneqq(v_{1},v_{2})$. Then, we have
@@ 3.2 Full order system
Here, $\tau_{s},\tau_{f}\in\mathbb{R}_{\geq 0}$ denote clocks that measure the elapsed time since the most recent slow and fast transmission, respectively, and $\kappa_{s},\kappa_{f}\in\mathbb{Z}_{\geq 0}$ are counters for the number of slow and fast transmissions. The state space is defined as $\mathbb{X}\coloneqq\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{u_{s}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}$, and the SPNCS is modeled by the hybrid system (gosate12) with hybrid input $u_{s}$ and $v$.
$\mathcal{H}:\left\{\begin{aligned} \dot{\xi}&=F(\xi,u_{s},\dot{u}_{s},\dot{v},\epsilon),\ \xi\in\mathcal{C}^{\epsilon},\\ {\xi}^{+}&\in G(\xi,v),\ \xi\in\mathcal{D}_{s}^{\epsilon}\cup\mathcal{D}_{f}^{\epsilon},\end{aligned}\right$. (11)
@@ 4 Stability analysis
By definition of $f_{es}$, there exist $A_{H_{s}}=M_{s}A_{21}^{s}$ and $L_{s}=M_{s}|A_{22}^{s}|\underline{a}_{W_{s}}^{-1}$, such that for all $\delta_{x}\in\mathbb{R}^{n_{x}},\kappa_{s}\in\mathbb{Z}_{\geq 0}$ and almost all ${e_{s}}\in\mathbb{R}^{n_{e_{s}}}$, the following inequality holds,
@@
Similarly, by definition of $g_{ef}$, there exist $L_{f}=M_{f}|a_{22}^{f}|\underline{a}_{W_{f}}^{-1}$ and $A_{H_{f}}=M_{f}A_{21}^{f}$ such that for all $x\in\mathbb{R}^{n_{x}},\kappa_{f}\in\mathbb{Z}_{\geq 0}$, and almost all ${e_{f}}\in\mathbb{R}^{n_{y_{f}}}$, the following inequality holds:
@@
Consider the system (11) satisfying Assumptions 1, and Condition 21 and 23. Let $L_{s}$ and $L_{f}$ come from (18) and (20), respectively, $\gamma_{f}$ and $\gamma_{s}$ come from Conditions 21 and 23, respectively. Moreover, let $\lambda_{s}$ and $\lambda_{f}$ come from Assumption 1, and $\lambda_{f}^{*}\in[\lambda_{f},1)$. Then, for any $\tau_{\text{miati}}^{s}\leq\tau_{\text{mati}}^{s}<T(L_{s},\gamma_{s},\lambda_{s})$ and $\tau_{\text{mati}}^{f}\leq\epsilon T(L_{f},\gamma_{f},\lambda_{f}^{*})$, the following holds:
There exist $k,a,\epsilon^{*}>0$ and $\gamma_{v_{1}},\gamma_{v_{2}},\gamma_{\dot{u}_{s}}\geq 0$ such that for all $\epsilon\in(0,\epsilon^{*}]$, any solution $\xi$ to $\mathcal{H}$ with hybrid inputs $u_{s}$ and $v$, where $u_{s}$ is absolutely continuous and $\dot{u}_{s},v\in\mathcal{L}_{\infty}$, satisfies a DISS property, i.e.,
@@ 5 Illustrative example
To verify the bound (25), we conducted simulations for the example presented above using $\epsilon=0.016$, $\tau_{\text{mati}}^{s}=0.13~\text{s}$, and $\tau_{\text{mati}}^{f}=7~\text{ms}\leq\epsilon^{*}T^{*}$, with initial conditions $x_{p}(0)=(0.1,0.1)$, $z_{p}(0)=(-0.2,-0.2)$, and all other states initialized at the origin. The results are shown in Figures 2 and 3, where two scenarios are considered: (i) a nonzero control input with no measurement noise, and (ii) zero control input subject to measurement noise.

### v4

@@ Abstract
thanks: This work was supported by the Australian Research Council under the Discovery Project DP200101303, the France Australia collaboration project IRP-ARS CNRS and the ANR COMMITS ANR-23-CE25-0005, and by ANID through the FONDECYT Initiation grant 11250715. E-mail corresponding author: weixuanw@student.unimelb.edu.au.
@@ 1 Introduction
This work provides two main contributions. First, it introduces a systematic observer design framework for SPNCS with linear plants via LMI. In particular, the proposed observer explicitly incorporates key network-induced effects, including sampling and scheduling protocols. By doing so, it enables the derivation of observer gains that make the observer robust against network-induced errors and thereby maximize the MATI estimates for both slow- and fast-varying signals. This contribution provides a unified and computationally tractable method for observer synthesis in multi-time-scale NCS, where network constraints play a critical role. Second, the paper establishes a robustness analysis framework for observer design in SPNCS with linear plants subject to measurement noise and plant input. Because noisy measurements are sampled and transmitted over the network, the communication process inherently attenuates high-frequency noise. Following the idea in scheres2024robustifying, by analysing the received noise (i.e., the noise after it passes through the network), we remove the dependence of the ultimate bound on the time derivative of the measurement noise, leading to a more practical and implementable robustness guarantee.
@@ 3.1 Change of coordinates
Define the estimation errors on the slow and fast states as $\delta_{x}\coloneqq x_{o}-x_{p}$ and $\delta_{z}\coloneqq z_{o}-z_{p}$, respectively. Let
$v\coloneqq(v_{1},v_{2}),\quad\hat{v}\coloneqq(\hat{v}_{1},\hat{v}_{2}).$ (8)
Then, we have
@@ 3.2 Full order system
Here, $\tau_{s},\tau_{f}\in\mathbb{R}_{\geq 0}$ denote clocks that measure the elapsed time since the most recent slow and fast transmission, respectively, and $\kappa_{s},\kappa_{f}\in\mathbb{Z}_{\geq 0}$ are counters for the number of slow and fast transmissions. The state space is defined as $\mathbb{X}\coloneqq\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{u_{s}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}$, and the SPNCS is modeled by the hybrid system (gosate12) with hybrid input $u_{s}$ and $v$, where $v$ is defined in (8).
$\mathcal{H}:\left\{\begin{aligned} \dot{\xi}&=F(\xi,\hat{v},u_{s},\dot{u}_{s},\dot{v},\epsilon),\ \xi\in\mathcal{C}^{\epsilon},\\ {\xi}^{+}&\in G(\xi,v),\ \xi\in\mathcal{D}_{s}^{\epsilon}\cup\mathcal{D}_{f}^{\epsilon},\end{aligned}\right$. (12)
@@ 4 Stability analysis
By definition of $f_{es}$ in (16), there exist $A_{H_{s}}=M_{s}A_{21}^{s}$ and $L_{s}=M_{s}|A_{22}^{s}|\underline{a}_{W_{s}}^{-1}$, such that for all $\delta_{x}\in\mathbb{R}^{n_{x}},\kappa_{s}\in\mathbb{Z}_{\geq 0}$ and almost all ${e_{s}}\in\mathbb{R}^{n_{e_{s}}}$, the following inequality holds,
@@
Similarly, by definition of $g_{ef}$ in (7), there exist $L_{f}=M_{f}|a_{22}^{f}|\underline{a}_{W_{f}}^{-1}$ and $A_{H_{f}}=M_{f}A_{21}^{f}$ such that for all $x\in\mathbb{R}^{n_{x}},\kappa_{f}\in\mathbb{Z}_{\geq 0}$, and almost all ${e_{f}}\in\mathbb{R}^{n_{y_{f}}}$, the following inequality holds:
@@
Consider the system (12) satisfying Assumptions 1, and Condition 22 and 24. Let $L_{s}$ and $L_{f}$ come from (19) and (21), respectively, $\gamma_{f}$ and $\gamma_{s}$ come from Conditions 22 and 24, respectively. Moreover, let $\lambda_{s}$ and $\lambda_{f}$ come from Assumption 1, and $\lambda_{f}^{*}\in[\lambda_{f},1)$. Then, for any $\tau_{\text{miati}}^{s}\leq\tau_{\text{mati}}^{s}<T(L_{s},\gamma_{s},\lambda_{s})$ and $\tau_{\text{mati}}^{f}\leq\epsilon T(L_{f},\gamma_{f},\lambda_{f}^{*})$, the following holds:
There exist $k,a,\epsilon^{*}>0$ and $\gamma_{v_{1}},\gamma_{v_{2}},\gamma_{\dot{u}_{s}}\geq 0$ such that for all $\epsilon\in(0,\epsilon^{*}]$, any solution $\xi$ to $\mathcal{H}$ with hybrid inputs $u_{s}$, $v_{1}$ and $v_{2}$, where $u_{s}$ is absolutely continuous and $\dot{u}_{s},v_{1},v_{2}\in\mathcal{L}_{\infty}$, satisfies a DISS property, i.e.,
@@ 5 Illustrative example
To verify the bound (26), we conducted simulations for the example presented above using $\epsilon=0.016$, $\tau_{\text{mati}}^{s}=0.15~\text{s}$, and $\tau_{\text{mati}}^{f}=7~\text{ms}\leq\epsilon^{*}T^{*}$, with initial conditions $x_{p}(0)=(0.1,0.1)$, $z_{p}(0)=(-0.2,-0.2)$, and all other states initialized at the origin. The results are shown in Figures 2 and 3, where two scenarios are considered: (i) a nonzero control input with no measurement noise, and (ii) zero control input subject to measurement noise.